Please help me answer the following questions given the data that I obtained, and please type your answer showing step by step calculations and formulas please

Experiment 2: Conservation of Heat Energy – Analysis and Discussion

Based on your experimental results, please answer the following questions:

  1. Given the initial and final volumes of the experiment, determine the initial mass of the water and the mass of the ice cube. Show your work. Record your answer in the table above.
  1. Given the initial and final temperatures, calculate the change in temperature of the water and ice after it melted. Show your work. Record your answer in the table above.
  1. Calculate the heat lost by the water and gained by the ice cube after it melted. Show your work. Record your answer in the table above.
  1. Use the conservation of energy to determine the amount of heat energy that it took to melt the ice. Show your work. Record your answer in the table above.
  1. It takes 334 J of heat energy to melt 1 g of ice (this is called the latent heat of fusion). Use this to predict the amount of energy it should take to melt your ice cube. Show your work. Record your answer in the table above.
  1. Compare this theoretical value to the value you calculated in Post-Lab Question 3 with a percent error calculation. Show your work. Record your answer in the table above.

Experimen Results/Observations Enter your data in the following tables and record your observations t 2: Conservation of Heat Ener Water Initial Final Water Volume and Temperature Data Volume (mL) 70 90 Temperature (C) 24.0 4.5 System Heat Data 70 20 Mass of water Mass of ice cube Change in water temperature (C Change in ice temperature (C Heat lost by water (J Head gained by ice after melting (J Heat used to melt ice (J) Theoretical heat to melt ice (J % Error
Answer #1

Mass of water = 70 ml = 70 g

mass of ice = final volume of water - initial volume water = 90 - 70 = 20 mL or 20 g

---------------------------------------------------------------------------------------------------------------------------

Change in water temperature = final - initial = 4.5 - 24 = -19.5 oC

--------------------------------------------------------------------------------------------------------------

Heat lost by water

Q = 70 g * 4.18 J / g oC * ( -19.5)

Q = - 5705.7 J ( the negative sign tells us that heat is lost)

--------------------------------------------------------------------------------------

Heat gained by ice. ( You did not mention initial temperature of ice)

here is the formula to find the heat gained by ice

Q1 = mass of ice * 4.18 * ( Tfinal - Tinitial )

after this, amount of heat required to melt the ice (Q2)

Q + Q1 + Q2 = 0

you have Q, Q1, solve for Q2